# LiD-FL: Towards List-Decodable Federated Learning

### v1

@@ Abstract
Federated learning is often used in environments with many unverified participants. Therefore, federated learning under adversarial attacks receives significant attention. This paper proposes an algorithmic framework for list-decodable federated learning, where a central server maintains a list of models, with at least one guaranteed to perform well. The framework has no strict restriction on the fraction of honest workers, extending the applicability of Byzantine federated learning to the scenario with more than half adversaries. Under proper assumptions on the loss function, we prove a convergence theorem for our method. Experimental results, including image classification tasks with both convex and non-convex losses, demonstrate that the proposed algorithm can withstand the malicious majority under various attacks.
@@ 1 Introduction
Although the federated learning paradigm improves privacy and the overall training efficiency, its training procedure is prone to malicious attacks (Karimireddy, He, and Jaggi 2022; Farhadkhani et al. 2024). Byzantine clients refers to arbitrarily behaving faulty clients except for changing the identities and behaviors of other clients. A variety of instantiations of the Byzantine attack model are designed, with the goal of preventing the model from converging or creating a backdoor to the system (Bagdasaryan et al. 2020; Shejwalkar et al. 2022). Previous study has shown that well-designed Byzantine attacks could degrade the global model or lead it to collapse (Blanchard et al. 2017; Shejwalkar and Houmansadr 2021; Baruch, Baruch, and Goldberg 2019; Xie, Koyejo, and Gupta 2020; Fang et al. 2020).
To address Byzantine attacks, a number of Byzantine-robust FL algorithms have been proposed. These algorithms usually leverage robust aggregation rules to filter out abnormal clients or control the influence of Byzantine clients to acceptable level (Blanchard et al. 2017; Pillutla, Kakade, and Harchaoui 2022a; Karimireddy, He, and Jaggi 2021; El Mhamdi, Guerraoui, and Rouault 2018; Allouah et al. 2023; Yi et al. 2024; Luan et al. 2024). The proposed Byzantine-robust FL algorithms are investigated from various perspectives, advancing both theoretical foundations and practical frontiers of this theme. It has been shown that previous solutions which use robust aggregators to train a global model have a breaking point of $1/2$ (Gupta and Vaidya 2020; Yin et al. 2018; Allouah et al. 2023; Farhadkhani et al. 2022; Allouah et al. 2024b). When the honest clients are less than one half, the robust aggregation rules lose efficacy.
@@
Although the emergence of FL has alleviated the data fragmentation issue to a certain degree (Konečný et al. 2016; McMahan et al. 2017), it is known to be vulnerable to device failure, communication breakdown and malicious attacks (El Mhamdi, Guerraoui, and Rouault 2018). In recent years, algorithmic frameworks are proposed to enhance the robustness of FL towards unreliable clients (Yin et al. 2018; Liu, Gupta, and Vaidya 2021; Allen-Zhu et al. 2021). Farhadkhani et al. propose a unified framework for Byzantine-robust distributed learning, RESAM, which utilizes resilient aggregation and distributed momentum to enhance the robustness of learning process.
@@
One of the key ingredients of prior robust FL algorithms is the Byzantine resilient aggregation rules, also called robust aggregators. The federated averaging algorithm (McMahan et al. 2017) uses naive average of local models to update the global model, which is effective in non-adversarial scenarios but could be easily destroyed by a single malicious client (Blanchard et al. 2017). Yin et al. propose two robust aggregations rules based on coordinate median and trimmed mean (TM), respectively. The theoretical analysis proves their statistical error rates for both convex and non-convex optimization objective, while their statistical rates for strongly convex losses are order-optimal (Yin et al. 2018). Geometric median (GM) has been proved to be an appropriate approximation of average as well as robust to adversarial clients (El-Mhamdi et al. 2023). The RFA algorithm leverages an approximate algorithm of GM to defend the federated learning against Byzantine attacks, which is computationally efficient and numerically stable (Pillutla, Kakade, and Harchaoui 2022a).
@@
Beside the statistical methods, many robust aggregators based on certain criteria of local updates are developed. The ARC (Allouah et al. 2024a), CClip (Karimireddy, He, and Jaggi 2021) and Norm (Sun et al. 2019) methods conduct norm prune on local model updates to resist Byzantine attacks, with a pre-computed of adaptive norm bound.
@@
The parameter server is assumed to hold no data in most of previous work about Byzantine-robust FL. The above algorithms, except for TM, do not demand prior knowledge about the fraction of Byzantine clients as input. However all of them require a Byzantine fraction of less than $1/2$ to ensure robustness (Gupta and Vaidya 2020; Yin et al. 2018). Limited work has been done to address the $51\%$ attack, where the Byzantine workers outnumber the honest workers. The FLTrust algorithm (Cao et al. 2021) assumes that the server is accessible to a tiny, clean dataset, which could be used to validate the local updates. Specifically, FLTrust weights local updates based on their cosine similarity to the model update on server dataset. However, FLTrust is inapplicable when public dataset is unavailable.
@@
The list-decoding technique was first introduced to the construction of error correcting codes and has then found applications in learning. A List-decoding model was introduced in (Balcan, Blum, and Vempala 2008) to solve the problem of finding proper properties of similarity functions for high-quality clustering. List-decodable learning is later proven to be an effective paradigm for the estimation, learning, and optimization tasks based on untrusted data (Charikar, Steinhardt, and Valiant 2017; Diakonikolas, Kane, and Stewart 2018). Karmalkar, Klivans, and Kothari propose a list-decodable learning algorithm for outlier-robust linear regression, with an arbitrary fraction of hostile outliers.
@@
These works show the usefulness of list-decodable learning for handling high-dimensional data and enhancing robust learning, even with significant corruption ratio. The skeleton of list-decodable learning is generating a list of $\mathrm{poly}(\frac{1}{\gamma})$ possible answers wherein at least one is correct, where $\gamma$ is the proportion of valid data. In this paper, we adapt the framework of list-decodable learning to Byzantine-robust federated learning.
@@ 2.1 Problem Setup
We consider a server-client federated learning system with one central server and $m$ clients $C=\{c_{1},\dots,c_{m}\}$. The server has no data and each client holds a block of data. The set of all training data is denoted by $X=\{x_{1},\dots,x_{n}\}$, and the indices of training data for each client $c_{j},j\in[m]$ is $D_{j}\subseteq[n]$. We consider the homogeneous setting in which the data points possessed by every client are independently and identically sampled from the same distribution $\mathcal{D}$.
For each training data point $x_{i}$, let $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ be its loss function. Given a model with parameter $w\in\mathbb{R}^{d}$, the loss of this model at $x_{i}$ is $f_{i}(w)$. We assume $f_{i}(w)$ is almost everywhere differentiable with respect to the parameter. Following the definition in (Allouah et al. 2024a), we define the local loss function for client $c_{j}$ as $f^{j}(w)=\frac{1}{|D_{j}|}\sum_{i\in D_{j}}f_{i}(w)$. The global loss function is the average of local loss functions $f(w)=\frac{1}{m}\sum_{j\in[m]}f^{j}(w)$. Then, our goal is to find the model parameter $w^{*}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{d}}f(w)$ to minimize the global loss.
@@ 2.2 Byzantine-Robust Federated Learning
Federated learning employs distributed stochastic gradient descent method (DSGD) to solve the above optimization problem. Specifically, each client holds a local model on its private dataset, and the server maintains a global model via aggregating the local models. The global model is optimized through stochastic gradient descent iteratively. For each global iteration $t\in\{0,\ldots,T-1\}$, the workflow of federated learning includes the following three steps:
(1) Broadcast: the central server samples a subset of clients $S_{t}\subseteq C$, and sends the current global model $w_{t}$ to them.
(2) Local Training: sampled clients train the model with local dataset privately, then return the local model updates to server synchronously. In detail, each sampled client $c_{j}\in S_{t}$ initially sets the local model as $w_{0}^{j}=w_{t}$, and updates it using a SGD optimizer for $\tau$ local batches. After local training, $c_{j}$ sends the local model update $u_{t}^{j}=w_{\tau}^{j}-w_{0}^{j}$ to the server.
(3) Global Update: the central server aggregates received local updates $U_{t}=\{u_{t}^{j}\}_{j\in S_{t}}$ via an aggregation rule $h:U\to\mathbb{R}^{d}$, e.g., taking the average, then updates the global model with aggregated model update $w_{t+1}=w_{t}+h(U_{t})$.
Now we look at a typical adversarial scenario where the server is reliable and $k$ out of $m$ clients are honest (or non-Byzantine) while the remainder are Byzantine (Lamport, Shostak, and Pease 2019). We use $\mathcal{H},\mathcal{B}\subseteq[m]$ to depict the indices of honest clients and Byzantine clients, respectively. Thus the fraction of honest clients is $\gamma=k/m$. The Byzantine clients are free to disregard the prescribed learning protocol and behave arbitrarily. In particular, they send malicious local updates to the server in the training process, trying to destroy the global model. Nonetheless, they are prohibited from corrupting other clients, altering the message of any other nodes, or hindering the transmission of messages between the server and any honest clients.
Therefore, in the Byzantine-robust federated learning, we define the global loss function as the average of local loss functions over all honest clients $f(w)=\frac{1}{|\mathcal{H}|}\sum_{j\in\mathcal{H}}f^{j}(w)$. Then, we aim to find the model parameter $w^{*}=\operatorname*{arg\,min}_{w}f(w)$ to minimize this global loss, with only a $\gamma$ fraction of reliable clients in the system (Karimireddy, He, and Jaggi 2022).
@@ 3 List-Decodable Federated Learning
In this Section, we propose a list-decodable federated learning framework named LiD-FL. It is a two-phase framework. The first phase is similar to the typical workflow of federated learning except that the server holds a list of global models. And the second phase conduct the list updating through a voting procedure. Specifically, LiD-FL differs from the conventional robust FL in Section 2 from three aspects: (i) LiD-FL maintains a list of global models $L$ instead of a single global model. In particular, LiD-FL preserves $q$ global models on the server and optimize them concurrently. (ii) LiD-FL replaces the aggregator with a random sampler to update the global model. Since existing aggregators can not promise attaining honest updates with more than $1/2$ fraction of Byzantine clients, we randomly sample one client for local training and global update. (iii) LiD-FL introduces a voting procedure to update the global list, which ensures that at least one model in the list is valid.
@@
Step 1. Server Broadcast: the central server randomly chooses a global model $w_{t}=L_{t}[i],i\in[q]$ from the list and sends it to a sampled client. We note that only one client $c_{j}\in C$ is sampled uniformly at random in each round.
Step 2. Local Training: the sampled client $c_{j}$ optimizes the global model using its local dataset, then sends the model update $u_{t}^{j}$ to the server. If $c_{j}$ is a Byzantine client, it sends arbitrary information to the server.
Step 3. Global Update: the central server calculates a new global model using the received update $w^{\prime}_{t}=w_{t}+u_{t}^{j}$, and adds the new model into the global list.
Step 4. Voting Procedure: The server broadcasts the global model list $L_{t}$ to all clients, and each client has to vote for one model in the list. After receiving votes from all clients, the server discards the model which receives the least number of votes. In the local process, each honest client accesses a validation oracle, which depends on a local validation set. Each honest client $c_{j},j\in\mathcal{H}$ evaluates the loss of all global models via the validation oracle and votes to the model with the minimum loss value. Byzantine clients vote arbitrarily.
Steps 1-4 are conducted repeatedly for $T$ global rounds, then LiD-FL outputs the global model list $L_{T}$. Although not all models in the final list are valid, the clients could deploy the model they consider to be acceptable, which has been verified using their local dataset. Algorithm 1 shows the pseudocode of Lid-FL.
@@
Initialization: For the server, the global model list $L_{0}$, the number of global iterations $T$; for each client $c_{j}$: the local training round $\tau$, batch size $b$, learning rate $\ell$, initial momentum $v_{0}^{j}=0$, momentum coefficient $\mu$.
1:  for $t=0,\ldots,T-1$ do
2:     Server randomly samples one client $c_{j}\in C$ and one global model $w_{t}\in L_{t}$, then sends $w_{t}$ to $c_{j}$.
@@
4:        for $r=0,\ldots,\tau-1$ do
5:           Client $c_{j}$ computes a stochastic gradient on a local batch $g_{r}^{j}=\nabla f_{b}(w_{r}^{j})$, where $w_{0}^{j}=w_{t}$.
6:           Update the local momentum:
$v_{r+1}^{j}=\mu\cdot v_{r}^{j}+g_{r}^{j};$
$w_{r+1}^{j}=w_{r}^{j}-\ell\cdot v_{r+1}^{j}.$
7:        end for
8:        $u_{t}^{j}=w_{\tau}^{j}-w_{0}^{j}$.
9:        $c_{j}$ sends $u_{t}^{j}$ to the server.
10:     end if
11:     (A Byzantine client $c_{j}$ sends any arbitrary update $\tilde{u}_{t}^{j}$ to the server.)
12:     Server extends the global model list:
$w^{\prime}_{t}=w_{t}+u_{t}^{j};$
$L_{t}\leftarrow L_{t}\cup\{w^{\prime}_{t}\}.$
13:     Server broadcasts $L_{t}$ to all clients.
14:     for each honest client $c_{j},j\in\mathcal{H}$, in parallel do
15:        Evaluate the loss of all models and send the index of model with lowest loss to the server.
16:     end for
17:     (A Byzantine client sends an arbitrary index to the server.)
18:     Server eliminates the model which wins least vote $\hat{w}_{t}$ to update the list:
@@
In the voting phase, each client has to cast one vote among the $(q+1)$ candidate models. Then the top $q$ candidates with the highest number of votes are preserved by the server. To ensure that at least one model voted by honest clients is maintained, the size of the global model list $q$ must satisfy $q\geq\lfloor 1/\gamma\rfloor$. In conclusion, to guarantee the reliability of the global model list, we set $q\geq\lfloor 1/\gamma\rfloor$. We compare the efficacy of LiD-FL with different values of $q$ in Appendix B.5, and the results show that upon satisfying the bound, a larger list size attains better stability while keeping the high performance.
@@
We use a simple random sampler to obtain local updates, which is valid and computationally efficient. However, the update of only one client is collected in each global round, which does not utilize the parallel computation power of clients. A more efficient way is to use robust aggregation rules to generate a list of aggregated updates. A preliminary experimental result is given in Appendix B.6, which shows that by leveraging aggregation rules, the algorithm performs better.

### v3

@@ Abstract
Federated learning is often used in environments with many unverified participants. Therefore, federated learning under adversarial attacks receives significant attention. This paper proposes an algorithmic framework for list-decodable federated learning, where a central server maintains a list of models, with at least one guaranteed to perform well. The framework has no strict restriction on the fraction of honest clients, extending the applicability of Byzantine federated learning to the scenario with more than half adversaries. Assuming the variance of gradient noise in stochastic gradient descent is bounded, we prove a convergence theorem of our method for strongly convex and smooth losses. Experimental results, including image classification tasks with both convex and non-convex losses, demonstrate that the proposed algorithm can withstand the malicious majority under various attacks. Code: https://github.com/jerry907/LiD-List-Decodable-Federated-Learning.
@@ 1 Introduction
Although the federated learning paradigm improves privacy and the overall training efficiency, its training procedure is prone to malicious attacks (Karimireddy et al. 2022; Farhadkhani et al. 2024). Byzantine clients refers to arbitrarily behaving faulty clients except for changing the identities and behaviors of other clients. A variety of instantiations of the Byzantine attack model are designed, with the goal of preventing the model from converging or creating a backdoor to the system (Bagdasaryan et al. 2020; Shejwalkar et al. 2022). Previous study has shown that well-designed Byzantine attacks could degrade the global model or lead it to collapse (Blanchard et al. 2017; Shejwalkar and Houmansadr 2021; Baruch et al. 2019; Xie et al. 2020; Fang et al. 2020).
To address Byzantine attacks, a number of Byzantine-robust FL algorithms have been proposed. These algorithms usually leverage robust aggregation rules to filter out abnormal clients or control the influence of Byzantine clients to acceptable level (Blanchard et al. 2017; Pillutla et al. 2022a; Karimireddy et al. 2021; El Mhamdi et al. 2018; Allouah et al. 2023; Yi et al. 2024; Luan et al. 2024). The proposed Byzantine-robust FL algorithms are investigated from various perspectives, advancing both theoretical foundations and practical frontiers of this theme. It has been shown that previous solutions which use robust aggregators to train a global model have a breaking point of $1/2$ (Gupta and Vaidya 2020; Yin et al. 2018; Allouah et al. 2023; Farhadkhani et al. 2022; Allouah et al. 2024b). When the honest clients are less than one half, the robust aggregation rules lose efficacy.
@@
Although the emergence of FL has alleviated the data fragmentation issue to a certain degree (Konečný et al. 2016; McMahan et al. 2017), it is known to be vulnerable to device failure, communication breakdown and malicious attacks (El Mhamdi et al. 2018). In recent years, algorithmic frameworks are proposed to enhance the robustness of FL towards unreliable clients (Yin et al. 2018; Liu et al. 2021; Allen-Zhu et al. 2021). Farhadkhani et al. propose a unified framework for Byzantine-robust distributed learning, RESAM, which utilizes resilient aggregation and distributed momentum to enhance the robustness of learning process.
@@
One of the key ingredients of prior robust FL algorithms is the Byzantine resilient aggregation rules, also called robust aggregators. The federated averaging algorithm (McMahan et al. 2017) uses naive average of local models to update the global model, which is effective in non-adversarial scenarios but could be easily destroyed by a single malicious client (Blanchard et al. 2017). Yin et al. propose two robust aggregations rules based on coordinate median and trimmed mean (TM), respectively. The theoretical analysis proves their statistical error rates for both convex and non-convex optimization objective, while their statistical rates for strongly convex losses are order-optimal (Yin et al. 2018). Geometric median (GM) has been proved to be an appropriate approximation of average as well as robust to adversarial clients (El-Mhamdi et al. 2023). The RFA algorithm leverages an approximate algorithm of GM to defend the federated learning against Byzantine attacks, which is computationally efficient and numerically stable (Pillutla et al. 2022a).
@@
Beside the statistical methods, many robust aggregators based on certain criteria of local updates are developed. The ARC (Allouah et al. 2024a), CClip (Karimireddy et al. 2021) and Norm (Sun et al. 2019) methods conduct norm prune on local model updates to resist Byzantine attacks, with a pre-computed of adaptive norm bound.
@@
The parameter server is assumed to hold no data in most of previous work about Byzantine-robust FL. The above algorithms, except for TM, do not demand prior knowledge about the fraction of Byzantine clients as input. However all of them require a Byzantine fraction of less than $1/2$ to ensure robustness (Gupta and Vaidya 2020; Yin et al. 2018). Limited work has been done to address the $51\%$ attack, where the Byzantine clients outnumber the honest clients. The FLTrust algorithm (Cao et al. 2021) assumes that the server is accessible to a tiny, clean dataset, which could be used to validate the local updates. Specifically, FLTrust weights local updates based on their cosine similarity to the model update on server dataset. Although FLTrust could tolerate more than half Byzantine clients, it is inapplicable when public dataset is unavailable.
@@
The list-decoding technique was first introduced to the construction of error correcting codes and has then found applications in learning. A List-decoding model was introduced in (Balcan et al. 2008) to solve the problem of finding proper properties of similarity functions for high-quality clustering. List-decodable learning is later proven to be an effective paradigm for the estimation, learning, and optimization tasks based on untrusted data (Charikar et al. 2017; Diakonikolas et al. 2018). Karmalkar et al. propose a list-decodable learning algorithm for outlier-robust linear regression, with an arbitrary fraction of hostile outliers.
@@
These studies show the usefulness of list-decodable learning for handling high-dimensional data and enhancing robust learning, even with a significant corruption ratio. The key idea of list-decodable learning is to generate a list of $\mathrm{poly}(\frac{1}{\gamma})$ possible answers including at least one correct answer, where $\gamma$ is the proportion of valid data. In this paper, we adapt the framework of list-decodable learning to Byzantine-robust federated learning.
@@ 2.1 Problem Setup
We consider a server-client federated learning system with one central server and $m$ clients $C=\{c_{1},\dots,c_{m}\}$. The server has no data and each client holds a block of data. The set of all training data is denoted by $Z=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $(x_{i},y_{i})$ represents an input-output pair. The indices of training data for each client $c_{j},j\in[m]$ is $D_{j}\subseteq[n]$. We consider the homogeneous setting in which the data possessed by every client are independently and identically sampled from the same distribution $\mathcal{D}$.
For each input-output pair $(x_{i},y_{i})$, let $f_{i}:\mathbb{R}^{d}\times Z\to\mathbb{R}$ be its loss function. Given a model with parameter $w\in\mathbb{R}^{d}$, the loss of this model at $(x_{i},y_{i})$ is $f_{i}(w)$. We assume $f_{i}(w)$ is almost everywhere differentiable with respect to the parameter. Following the definition in (Allouah et al. 2024a), we define the local loss function for client $c_{j}$ as $f^{(j)}(w)=\frac{1}{|D_{j}|}\sum_{i\in D_{j}}f_{i}(w)$. The global loss function is the average of local loss functions $f(w)=\frac{1}{m}\sum_{j\in[m]}f^{(j)}(w)$. Then, our goal is to find the model parameter $w^{*}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname*{arg\,min}_{w\in\mathbb{R}^{d}}f(w)$ to minimize the global loss.
@@ 2.2 Byzantine-Robust Federated Learning
Federated learning employs distributed stochastic gradient descent method to solve the above optimization problem. Specifically, each client holds a local model on its private dataset, and the server maintains a global model via aggregating the local models. The global model is optimized through stochastic gradient descent (SGD) iteratively. For each global iteration $t\in\{0,\ldots,T-1\}$, the workflow of federated learning includes the following three steps:
(1) Server Broadcast: the server samples a subset of clients $S_{t}\subseteq C$ and sends the current global model $\tilde{w}_{t}$ to them.
(2) Local Training: Each sampled client $c_{j}\in S_{t}$ initially sets the local model $w_{0}^{(j)}$ as $\tilde{w}_{t}$, and optimizes it via SGD for $\tau$ local batches. Then $c_{j}$ sends the local model update $u_{t}^{(j)}=w_{\tau}^{(j)}-w_{0}^{(j)}$ to the server.
(3) Global Update: the server aggregates received local model updates $U_{t}=\{u_{t}^{(j)}\}_{j\in S_{t}}$ via an aggregation rule $h:\mathbb{R}^{d\times|S_{t}|}\to\mathbb{R}^{d}$, e.g., taking the average, then updates the global model with aggregated update $w_{t+1}=w_{t}+h(U_{t})$.
Now we look at a typical adversarial scenario where the server is reliable and $k$ out of $m$ clients are honest (or non-Byzantine) while the remainder are Byzantine (Lamport et al. 2019). We use $\mathcal{H},\mathcal{B}\subseteq[m]$ to depict the indices of honest clients and Byzantine clients, respectively. Thus the fraction of honest clients is $\gamma=k/m$. The Byzantine clients are free to disregard the prescribed learning protocol and behave arbitrarily. In particular, they send malicious local model updates to the server in the training process, trying to destroy the global model. Nonetheless, they are prohibited from corrupting other clients, altering the message of any other nodes, or hindering the transmission of messages between the server and any honest clients.
Therefore, in the Byzantine-robust federated learning, we define the global loss function as the average of local loss functions over all honest clients $f(w)=\frac{1}{|\mathcal{H}|}\sum_{j\in\mathcal{H}}f^{(j)}(w)$. Then, we aim to find the model parameter $w^{*}=\operatorname*{arg\,min}_{w}f(w)$ to minimize this global loss, with only a $\gamma$ fraction of reliable clients in the system (Karimireddy et al. 2022).
@@ 3 List-Decodable Federated Learning
In this section, we propose a list-decodable federated learning framework named LiD-FL. It is a two-phase framework. The first phase is similar to the typical workflow of federated learning except that the server holds a list of global models. And the second phase conducts the list updating through a voting procedure. Specifically, LiD-FL differs from the conventional robust FL in Section 2 from three aspects: (i) LiD-FL maintains a list of global models $L$ instead of a single global model. In particular, LiD-FL preserves $q$ global models on the server and optimizes them concurrently. (ii) LiD-FL replaces the aggregator with a random sampler to update the global model. Since existing aggregators can not promise attaining honest updates with more than $1/2$ fraction of Byzantine clients, we randomly sample one client for local training and global update. (iii) LiD-FL introduces a voting procedure to update the global model list, which ensures that at least one model in the list is valid.
@@
Step 1. Server Broadcast: the server randomly chooses a global model $\tilde{w}_{t}$ from the current list $L_{t}$ and sends it to a sampled client. We note that only one client $c_{j}\in C$ is sampled uniformly at random in each round.
Step 2. Local Training: the sampled client $c_{j}$ optimizes the global model using its local dataset, then sends the model update $u_{t}^{(j)}$ to the server. If $c_{j}$ is a Byzantine client, it sends arbitrary information to the server [^1] If a Byzantine client does not respond, the server can simply ignore it or fake an arbitrary update for it. The same strategy can be applied to the voting step..
Step 3. Global Update: the server calculates a new global model using the received update $w^{\prime}_{t}=w_{t}+u_{t}^{(j)}$, and adds the new model into the global model list.
Step 4. Local Voting: The server broadcasts the list $L_{t}$ to all clients. Every client votes for one model in $L_{t}$. Each honest client $c_{j},j\in\mathcal{H}$ evaluates the loss of all global models via a validation oracle. It then votes for the model with the minimum loss value. Byzantine clients vote arbitrarily.
Step 5. List Update: The server discards the model which receives the least number of votes to obtain a new list $L_{t+1}$
Steps 1-5 are conducted repeatedly for $T$ global rounds, then LiD-FL outputs the global model list $L_{T}$. Although not all models in the final list are valid, the clients could deploy the model they consider to be acceptable by evaluating models in the list using their local dataset. Algorithm 1 shows the pseudocode of LiD-FL.
@@
Initialization: For the server, the global model list $L_{0}$, the number of global iterations $T$; for each client $c_{j}$: the number of local training rounds $\tau$, batch size $b$, learning rate $\ell$.
1:  for $t\in\{0,\ldots,T-1\}$ do
2:     Server randomly samples one client $c_{j}\in C$ and one global model $\tilde{w}_{t}\in L_{t}$, then sends $\tilde{w}_{t}$ to $c_{j}$.
@@
4:        $w_{0}^{(j)}\leftarrow\tilde{w}_{t}$.
5:        for $r\in\{0,\ldots,\tau-1\}$ do
6:           Client $c_{j}$ computes a stochastic gradient $g_{r}^{(j)}$ on a batch of local data.
7:           Update the local model:
$w_{r+1}^{(j)}\leftarrow w_{r}^{(j)}-\ell\cdot g_{r}^{(j)}.$
8:        end for
9:        $u_{t}^{(j)}\leftarrow w_{\tau}^{(j)}-w_{0}^{(j)}$.
10:        $c_{j}$ sends $u_{t}^{(j)}$ to the server.
11:     end if
12:     (A Byzantine client $c_{j}$ sends arbitrary update $\tilde{u}_{t}^{(j)}$ to the server.)
13:     Server extends the global model list:
$\tilde{w}^{\prime}_{t}\leftarrow\tilde{w}_{t}+u_{t}^{(j)},$
$L_{t}\leftarrow L_{t}\cup\{\tilde{w}^{\prime}_{t}\}.$
14:     Server broadcasts $L_{t}$ to all clients.
15:     for each honest client $c_{j},j\in\mathcal{H}$, in parallel do
16:        Evaluate the lo ss of all models using its validation set, and send the index of the model with lowest loss to the server.
17:     end for
18:     (A Byzantine client sends an arbitrary index to the server.)
19:     Server removes the model $\hat{w}_{t}$ which receives the least number of votes from the list:
@@
In the voting procedure, each client casts one vote among the $(q+1)$ candidate models. Then the top $q$ candidates with the highest number of votes are preserved by the server. To ensure that at least one model voted by honest clients is maintained, the size of the global model list $q$ must satisfy $q\geq\lfloor 1/\gamma\rfloor$. We compare the efficacy of LiD-FL with different values of $q$ in the appendix, and the results show that when the bound is satisfied, a larger list size attains better stability while maintaining comparable accuracy.
@@
We use a simple random sampler to obtain local updates, which is valid and computationally efficient. However, the update of only one client is collected in each global round, which does not utilize the parallel computation power of clients. A more efficient way is to use robust aggregation rules to generate a list of aggregated updates. A preliminary experimental result is given in the appendix, which shows that by leveraging aggregation rules, the algorithm performs better.